a) Points A and B are at 0.05m and 0.03m from a charge +13 x10-6C respectively. Determine the required work against the direction of the electric field to displace a charge +15 x10-6C from point A to point B.

b) A charge q1=+8 x10-6C is at the origin (x=0mm), and other charge q2=-3 x10-6C is at x=60mm. Find the work done by the electric field if a charge q3=-4 x10-6C is displaced from point x=15mm to point x=5mm.



1
Expert's answer
2021-10-26T17:17:53-0400

a)


"W=(9\cdot10^{9})\frac{(15\cdot10^{-6})(13\cdot10^{-6})}{0.03}-\\(9\cdot10^{9})\frac{(15\cdot10^{-6})(13\cdot10^{-6})}{0.05}=23.4\ J"

b)


"E_1=(9\cdot10^{9})\frac{(3\cdot10^{-6})(4\cdot10^{-6})}{0.045}-\\(9\cdot10^{9})\frac{(8\cdot10^{-6})(4\cdot10^{-6})}{0.015}=-16.8\ J"

"E_2=(9\cdot10^{9})\frac{(3\cdot10^{-6})(4\cdot10^{-6})}{0.055}-\\(9\cdot10^{9})\frac{(8\cdot10^{-6})(4\cdot10^{-6})}{0.005}=-55.6\ J\\W=-55.6+16.8=-38.8\ J"